Two charges are placed on the x-axis. A charge q1= 6µC is placed at x=3cm and charge

q2=-12µC charge is placed at x=5cm. a) Determine the magnitude and direction of the force

on the q1. b) Determine the magnitude and direction of the electric field at the origin.


1
Expert's answer
2021-04-26T17:12:28-0400

a)


"F=(9\cdot10^{9})\frac{(6\cdot10^{-6})(12\cdot10^{-6})}{(0.05-0.03)^2}=1620\ N"

Positive x-direction.

b)


"E=-(9\cdot10^{9})\frac{(6\cdot10^{-6})}{(0.05-0.03)^2}+(9\cdot10^{9})\frac{(12\cdot10^{-6})}{(0.05-0.03)^2}\\=1.35\cdot10^8\frac{N}{C}"

Positive x-direction.